Assume that 12 jurors are randomly selected from a population in which 72% of the people are Mexican-Americans. Refer to the probability distribution table below and find the indicated probabilities.
x
   P(x)
0   0+
1   0+
2   0.0001
3   0.0009
4   0.005
5   0.0207
6   0.062
7   0.1367
8   0.2197
9   0.2511
10   0.1937
11   0.0906
12   0.0194


Find the probability of exactly 6 Mexican-Americans among 12 jurors.
P(x=6)=

Find the probability of 6 or fewer Mexican-Americans among 12 jurors.
P(x≤6)=

Does 6 Mexican-Americans among 12 jurors suggest that the selection process discriminates against Mexican-Americans?

no
yes
Answer #1

From the given table:

P(Exactly 6) = P(X = 6) = 0.062

P(x < 6) = P(0) + P(1) .... + P(6) = 0 + 0 + 0.0001 + 0.0009 + 0.005 + 0.0207 + 0.062 = 0.0887

Since probability of 6 or less is greater than 0.05, the answer is No.